OpenStudy (anonymous):

Suppose you roll 2 dice. Find the probability of rolling a sum of 9

OpenStudy (anonymous):

@haleyelizabeth2017

OpenStudy (anonymous):

I know there are 4 possible outcomes:6+3, 5+4, 4+5, 3+6, so wouldn't it be 1/4?

OpenStudy (anonymous):

possible

OpenStudy (anonymous):

or 4/12

OpenStudy (anonymous):

because there are 4 outcomes and 12 numbers all together on the dice

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

Here the total number of points in event space is 6^2=36 and 9=3+6=6+3 =5+4=4+5 Hence only 4 points corresponds to the required solution required probability is 4/36=1/9

sammixboo (sammixboo):

Okie dokie. If you don't understand something below then just tell me! :D Since you have two dice, and there is 4 outcomes(6+3, 5+4, 4+5, and 3+6), and 36 possible outcomes, so we have \(\sf \large \color{magenta}{\frac{4}{36}}\) and if you simplify that you get \(\sf \large \color{magenta}{\frac{1}{9}}\)

sammixboo (sammixboo):

Sorry for taking forever o _ o My internet cut off for a bit

OpenStudy (anonymous):

Thank You Guys =D

sammixboo (sammixboo):

No problem =)

OpenStudy (anonymous):

No prob sammi, n ok

sammixboo (sammixboo):

If you need anymore help just tag me in your questions :P

OpenStudy (anonymous):

ok